What mass of O2 is needed to increase the temperature of 35 grams polyethylene from 250K to 375. K? (specific heat of polyethylene is 0.551 cal/gK)
P4 + 5O2 --> P4O10 ∆H = - 3013 kJ/mol
1
Expert's answer
2020-12-10T07:47:09-0500

First calculate the amount of heat needed to increase the temperature of 35 grams of polyethylene from 250K to 375K:

"Q = cm\Delta T = 0.551 * 35 * (375-250) = 2410.6 (cal)"

Convert calories to joules (1 cal = 4.184 J):

"Q = 2410.6 * 4.184 = 10086 (J) = 10.09 (kJ)"

Find the mass of O2 needed to produce 3013 kJ of heat (5 moles according to the equation):

"m(O_2) = M*\nu = 32 * 5 = 160 (g)"

If 160 g of Oare needed to produce 3013 kJ of heat, then calculate the amount of NO2 needed to produce 10.09 kJ of heat:

"m(O_2) = 160 * 10.09\/3013 = 0.536 (g)"

Answer: 0.536 g of O2 is needed to increase the temperature of 35 grams of polyethylene from 250K to 375K.